OpenStudy (migitmack):

Fantastic Furniture builds sofas. It costs $2,725 to build 9 sofas and $3,850 to build 14 sofas. Which equation represents the cost, C(x), as a linear function of the number of sofas, x? C(x) = 225x - 700 C(x) = 225x + 700 C(x) = 700x - 225 C(x) = 700x + 225

OpenStudy (migitmack):

PLz help me

OpenStudy (anonymous):

Ok I get it. Basically the x number will be 9 & 14. whenever you put both numbers into each of the equations one at a time, you should get $2,725 & $3,850. Do you get that?

OpenStudy (anonymous):

Once we do this the only one that works is the 2nd equation. 225(9)+700= 2025+700=2725 So this is correct. 225(14)+700= 3150+700=3850 So this is also correct. This means our answer is B. :D

OpenStudy (migitmack):

OK thx i get it now

OpenStudy (anonymous):

No prob. and hey don't forget to give me a medal :)